OpenStudy (anonymous):

what is 1/2 to the 3rd power in fraction form?

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

\[ \frac{1}{8} \]

OpenStudy (anonymous):

1/8 because 2*2=4*2=8 (:

OpenStudy (anonymous):